cosq = 5/12 and tanq is negative, find the exact value for cscq.


Answer by AnlytcPhil(1810)

cosq = 5/12 and tanq is negative, find the exact value for cscq.

5/12 is a positive number and the cosine is positive in I and IV,
the tangent is negative in II and IV.  So, q can only be in IV.

Draw a picture          |
                        |
                        | x
                 -------|---------
                        |\   |
                        | \  |y
                        | r\ |
                        |   \|

Since cosq = 5/12 = x/r, put 5 for the x and 12 for the r

Draw a picture          |
                        |
                        | 5
                 -------|---------
                        |\   |
                        | \  |y
                        |12\ |
                        |   \|

Use the Pythagorean theorem 

                     x² + y² = r²
                     5² + y² = 12²
                     25 + y² = 144
                          y² = 119
                                 ___  
                           y = ±Ö119

Since y goes down, we select the negative sign.
                                 ___  
                           y = -Ö119

You are asked to find the cscq.
                   ___         ___
cscq = r/y = 12/(-Ö119) = -12/Ö119

Edwin
AnlytcPhil@aol.com